OpenStudy (anonymous):

Brianna is graphing the function f(x) = x2 + 6x + 5. What x-intercepts should Brianna use to graph f(x)? –5 and –1 –5 and 1 –1 and 5 1 and 5

OpenStudy (anonymous):

@jim_thompson5910 @AlexandervonHumboldt2 @triciaal

jimthompson5910 (jim_thompson5910):

an x-intercept is an x value that makes the whole right side equal to 0

OpenStudy (alexandervonhumboldt2):

so substtute them to the eqaution

OpenStudy (anonymous):

like How

jimthompson5910 (jim_thompson5910):

let's say someone claims that x = 1 is a root of f(x)

OpenStudy (alexandervonhumboldt2):

A

jimthompson5910 (jim_thompson5910):

that would mean they claim that f(1) = 0 let's find out f(x) = x^2 + 6x + 5 f(1) = (1)^2 + 6(1) + 5 ... replace every x with 1 f(1) = 1 + 6 + 5 f(1) = 12 the input was x = 1 and the output was y = 12 The output is NOT y = 0, so x = 1 is NOT a root or x-intercept

OpenStudy (triciaal):

what are the factors of +5 that add up to +6 1 and 5 so (x+1)(x+5) =0 x = -1 and x = -5